OpenStudy (anonymous):

how many tiles can she order if her cost is 7370.3??

OpenStudy (anonymous):

HELP ME PLZZZZZZZZZZZZ

OpenStudy (anonymous):

What is the cost per tile?

OpenStudy (anonymous):

$15.55

OpenStudy (anonymous):

(7370.3 dollars) / (15.55 dollars / tile) = answer in # of tiles.

OpenStudy (anonymous):

can u explain a little more plzzzz

OpenStudy (anonymous):

There will be dollars (the word dollars) in the numerator and the denominator, 'dollars will cancel out. "Tiles" is in the "denominator of the denominator" so it will flip up into the numerator. You are left with (7370.3 / 15.55) "tiles". Once you do the fraction, there is no more numerator/denominator, so your answer is in tiles.

OpenStudy (anonymous):

oh i get it thanks u

OpenStudy (anonymous):

The hard part is how to set the problem up.

OpenStudy (anonymous):

A great trick is to get your answer unit (tiles) in either the numerator or if you are dividing by something with "tiles" in it (dollars per tile), make sure the divisor has tiles in the denominator. It will flip up into the numerator.

OpenStudy (anonymous):

I hope you read this trick, it is very helpful. Did you?

OpenStudy (anonymous):

alittle

jimthompson5910 (jim_thompson5910):

Let x = number of tiles if one tile costs $15.55, then two cost 15.55*2 = 31.10 dollars, three cost 15.55*3 = 46.65 dollars, etc In general, x tiles cost a total of 15.55x dollars If the total cost is 7370.3, then we know that 15.55x = 7370.3 Solve that for x to find your answer. Remember to round down (if you get a decimal answer)

OpenStudy (anonymous):

thank you evry much

jimthompson5910 (jim_thompson5910):

np